# A weak regularity lemma for polynomials

### v1

@@ 3.1 Properties of univariate degree
    Fix all but the $i$-ith variable to constants so that the resulting univariate polynomial is nonzero. (It is indeed clear that such an assignment exists: we may assume that all variables in $P$ have degree smaller than $|\mathbb{F}|$. Then, $P=\sum_{j}x_{i}^{j}C_{j}$, with some $C_{j}$ a nonzero polynomial and thus also a nonzero function on $\mathbb{F}^{n-1}$).
@@ 3.2 Proofs of the applications
Put $\mathbb{F}=\mathbb{F}_{q}$. Assume $\mathbf{P}$ is nonzero, as otherwise $\operatorname{rk}_{t}(0)=0$ and we are done. Put $\epsilon=1-d/q$. By Theorem 2.2, $\mathbf{P}$ has a weak $\epsilon$-regular decomposition $\mathbf{P}=\mathbf{F}(\mathbf{X})$, with $\mathbf{X}\in\operatorname{Form}^{r}$, of size $r\leq(2m)^{2^{d(1+o(1))}}$. We assume without loss of generality that each decomposition $P_{i}=F_{i}(\mathbf{X})$ is irredundant (by removing monomials from the support of $F_{j}$ if necessary). Since $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]\setminus\mathbb{F}[\mathbf{X}^{\prime}]$, there is a component $P_{j}$ of $\mathbf{P}$ such that $P_{j}\in\mathbb{F}[\mathbf{X}]\setminus\mathbb{F}[\mathbf{X}^{\prime}]$, and therefore $\deg_{1}(F_{j})>0$. Apply Corollary 3.13 on $P_{j}\in\operatorname{Poly}_{d}(\mathbb{F})$ to obtain a line $\ell\subseteq\mathbb{F}^{r}$ in direction $e_{1}$ such that $|F_{j}(\ell)|\neq 1$ and $\Pr[\mathbf{X}\in\ell]>(1-d/q)q^{-(r-1)}$. Now apply Lemma 3.10 on $\ell$, using the weak $\epsilon$-regularity of $\mathbf{X}$ and our choice of $\epsilon$, to deduce that
@@
$\operatorname{rk}_{d/u}(\mathbf{P})\leq\sum_{i=0}^{d}r^{i}\leq 2r^{d}\leq(2m)^{2^{d(1+o(1))}}.\qed$
Theorem 1.2 follows easily from Theorem 3.7. To see this, note that the latter gives for any $\mathbf{P}\in\operatorname{Poly}_{d}^{m}$ with $\operatorname{udeg}(\mathbf{P})=u$ a $\sum^{[r]}\prod^{[u]}\sum\prod^{[d/u]}$-formula with top fan-in $r\leq f(m,d)$ (using our definition of $\operatorname{rk}_{t}(\mathbf{P})$ for a polynomial tuple $\mathbf{P}$, extending Definition 3.1). This implies a similar formula for any $P\in\operatorname{Poly}_{d}$ with $P\in\mathbb{F}[\mathbf{P}]$, as needed in Theorem 1.2.

### v2

@@ 3.1 Properties of univariate degree
    Fix all but the $i$-ith variable to constants so that the resulting univariate polynomial is nonzero. (It is indeed clear that such an assignment exists: we may assume that all variables in $P$ have degree smaller than $|\mathbb{F}|$. Then, $P=\sum_{j}x_{i}^{j}C_{j}$, with some $C_{j}$ a nonzero polynomial and thus also a nonzero function on $\mathbb{F}^{n-1}$.)
@@ 3.2 Proofs of the applications
Put $\mathbb{F}=\mathbb{F}_{q}$. Assume $\mathbf{P}$ is nonzero, as otherwise $\operatorname{rk}_{t}(0)=0$ and we are done. Put $\epsilon=1-d/q$. By Theorem 2.2, $\mathbf{P}$ has a weak $\epsilon$-regular decomposition $\mathbf{P}=\mathbf{F}(\mathbf{X})$, with $\mathbf{X}\in\operatorname{Form}^{r}$, of size $r\leq(2m)^{2^{d(1+o(1))}}$. We will prove that $\deg(\mathbf{X})\leq\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P})$.
We assume without loss of generality that each decomposition $P_{i}=F_{i}(\mathbf{X})$ is irredundant (by removing monomials from the support of $F_{j}$ if necessary). Since $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]\setminus\mathbb{F}[\mathbf{X}^{\prime}]$, there is a component $P_{j}$ of $\mathbf{P}$ such that $P_{j}\in\mathbb{F}[\mathbf{X}]\setminus\mathbb{F}[\mathbf{X}^{\prime}]$, and therefore $\deg_{1}(F_{j})>0$. Apply Corollary 3.13 on $P_{j}\in\operatorname{Poly}_{d}(\mathbb{F})$ to obtain a line $\ell\subseteq\mathbb{F}^{r}$ in direction $e_{1}$ such that $|F_{j}(\ell)|\neq 1$ and $\Pr[\mathbf{X}\in\ell]>(1-d/q)q^{-(r-1)}$. Now apply Lemma 3.10 on $\ell$, using the weak $\epsilon$-regularity of $\mathbf{X}$ and our choice of $\epsilon$, to deduce that
@@
$\operatorname{rk}_{d/u}(\mathbf{P})\leq\binom{r+d}{d}=\sum_{i=0}^{d}\binom{r+i-1}{i}\leq\sum_{i=0}^{d}r^{i}\leq(2m)^{2^{d(1+o(1))}}.\qed$
Theorem 1.2, like Theorem 3.7, follows from the weak regularity lemma.
Proof of Theorem 1.2.
Let $P\subseteq\mathbb{F}[\mathbf{P}]$ with $\operatorname{udeg}(\mathbf{P})=u$. As in the proof of Theorem 3.7, using the weak regularity lemma we obtain a decomposition $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ with $\mathbf{X}\in\operatorname{Form}^{s}$ of size $s\leq(2m)^{2^{d(1+o(1))}}$ such that $\deg(\mathbf{X})\leq d/u$. Observe that $P\in\mathbb{F}[\mathbf{P}]\subseteq\mathbb{F}[\mathbf{X}]$. Using Lemma 3.11, it follows that we can write
$P=\sum_{i=1}^{r}c_{i}\prod_{j}Q_{i,j}.$ (8)
with $Q_{i,j}\in\mathbf{X}$, $c_{i}\in\mathbb{F}$, $\deg(\prod_{j}Q_{i,j})\leq d$, and $r\leq\sum_{i=0}^{d}s^{i}\leq(2m)^{2^{d(1+o(1))}}$. Note that $\deg(Q_{i,j})\leq\deg(\mathbf{X})\leq d/u$.
We claim that we can guarantee that the number of multiplicands in each product $\prod_{j}Q_{i,j}$ is at most $2u$. Note that we can express each product as a product of homogeneous polynomials $\prod_{j=1}^{a_{i}}Q^{\prime}_{i,j}$ all of whose degrees, except at most one of them, satisfy $\frac{1}{2}(d/u)<\deg(Q^{\prime}_{i,j})\leq d/u$; indeed, this follows by repeatedly using the fact that if $\deg(Q_{i,j_{1}}),\deg(Q_{i,j_{2}})\leq\frac{1}{2}(d/u)$ then $\deg(Q_{i,j_{1}}Q_{i,j_{2}})\leq d/u$. Thus, $d\geq\sum_{j=1}^{a_{i}}\deg(Q^{\prime}_{i,j})>(a_{i}-1)\frac{1}{2}(d/u)$, so $a_{i}\leq 2u$ as claimed. Thus, (8) gives a $\sum^{[r]}\prod^{[2u]}\sum\prod^{[d/u]}$-formula for $P$, as needed.
Finally, if $P$ is homogeneous then, since every member of $\mathbf{X}$ is homogeneous, without loss of generality the products in the formula satisfy $\deg(\prod_{j}Q_{i,j})=\deg(P)$ for every $i$, and therefore the formula is homogeneous, as desired. ∎